QUESTION 21 What is the pH of a solution with a hydronium (Ho concentration of 2.7 x 10-PM.QUESTION 22 What is the hydronium (H307) concentration of a solution with a pH of 4.03?QUESTION 23 What is the hydroxide (OH) concentration of a solution with a hyaronium (H307) concentration of 5 x 10-8?
Answer #1

21.

pH = - log [H3O+]

So, pH = - log (2.7×10-3)

= 2.57.

22.

pH = - log [H3O+]

Given, pH = 4.08

So, - log [H3O+] = 4.08

Or, [H3O+] = 10-4.08

Or, [H3O+] = 8.3×10-5 M.

23.

At 250 c

[H3O+] [OH-] = 10-14

Or, 5×10-8 × [OH-] = 10-14

Or, [OH-] = (10-14 /5×10-8)

Or, [OH-] = 2.0× 10-7 M.